OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to 6= -4x- 2passes through (-16,-11).

OpenStudy (anonymous):

So first we should find the slope. Do you know what that is, if the equation is parallel?

OpenStudy (anonymous):

actually it is suppose to be perpendicular (sorry little typo)

OpenStudy (anonymous):

If slope- intercept form is y=mx+b and m=the slope. What is it? Perpendicular if fine!

OpenStudy (anonymous):

the slope will be [1/4\]

OpenStudy (jdoe0001):

is that 6= -4x- 2y?

OpenStudy (jdoe0001):

the "y" is conspicuous for its absence

OpenStudy (anonymous):

You're right my bad!

OpenStudy (jdoe0001):

so chrissychris, what's the equation?

OpenStudy (anonymous):

What is the starting equation?

OpenStudy (anonymous):

no sorry let me re write the equation y=-4x-2

OpenStudy (anonymous):

That's fine so if we're trying to find the perpendicular equation for the point (-16, -11) the slope is still 1/4. So we insert 1/4 into the equation y=1/4x+b Next we must find b by inserting x & y which is (-16, -11)

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

-11=1/4(-16) + b -11=-4 +b Add 4 to both sides -8=b

OpenStudy (anonymous):

1/4x-8 right?

OpenStudy (anonymous):

Sorry -11+4= -7 so it would be y=1/4-7

OpenStudy (anonymous):

My bad. Check it so -11=1/4(-16)-7 Yes. It works.

OpenStudy (anonymous):

oh thank you so much

OpenStudy (anonymous):

No problem, any time! I hope you understood!